Use the following contribution margin statement for 20X9:

20X9
Sales volume (#units) 100
Revenue $7,500
  Variable costs $4,500
Contribution margin $3,000
  Fixed costs $1,800
Profit $1,200

e) How much is the breakeven volume?

Breakeven revenue?


f) How much is the margin of safety percentage (at current sales volume of 100 units)?
(enter percentages as a fraction of 1, i.e., enter 23.47% as 0.2347)

g) Based on the margin of safety computed in (f), will you start making a loss if sales drop by 30%?
(enter 1 for yes, 2 for no)

h) How much is the operating leverage (at current sales volume of 100 units)?
(enter percentages as a fraction of 1, i.e., enter 23.47% as 0.2347)

If fixed costs increase, will it increase or decrease the operating risk?
(enter 1=increase, 2=decrease)
Answer #1

Unit contribution Margin = Total contribution margin/Number of units

= 3000/100

= $30 per unit

Break even volume = Fixed costs/Unit contribution margin

= 1800/30

= 60 units

f)Margin of Safety % = (Sales – Break even sales)/Sales

= (100-60)/100

= 0.40

g)No, i.e. 2

since sales can drop by margin of safety i.e. 40% before incurring a loss

h)Operating leverage = Contribution margin/Net operating income

= 3000/1200

= 2.5 times

The operating risk will increase i.e. 1